# The Scale for the Expansion of the Universe: From Local Structures to Cosmology

### v1

@@ III Test Particle Geodesic
We can solve the motion analytically only for the Newtonian case, with the eccentricity parameter $e$ and the phase parameter $\eta$. These solutions describe the evolution of the radial coordinate $r$, the radial velocity $v_{r}$, and the tangential velocity $v_{t}$ as functions of time $t$. For bounded systems, where $e\in[0,1]$ the radial coordinate $r$ and time $t$ are given by Sandage (1986); Karachentsev et al. (2009):
$r=\frac{r_{\text{ta}}}{2}\left(1-e\cos\eta\right)\;,\quad t=\frac{t_{\text{ta}}}{\pi}\left(\eta-e\sin\eta\right)\;,$
$v_{r}=\frac{\pi}{2}\frac{r_{\text{ta}}}{t_{\text{ta}}}\frac{e\sin\eta}{1-e\cos\eta}\;,\quad v_{t}=\frac{\pi}{2}\frac{r_{\text{ta}}}{t_{\text{ta}}}\frac{\sqrt{1-e^{2}}}{1-e\cos\eta}\;,$ (7)
@@
where $r_{\text{ta}}$ and $t_{\text{ta}}$ are the turnaround radius and turnaround time, respectively, with $e\in[1,\infty]$. For the unbounded dwarf galaxies, the radial coordinate $r$ and time $t$ are expressed as:
$r=\frac{r_{\text{ta}}}{2}\left(e\cosh\eta-1\right)\;,\quad t=\frac{t_{\text{ta}}}{\pi}\left(e\sinh\eta-\eta\right)\;,$
@@
The enclosed mass is given by: $M=\pi^{2}r_{\text{ta}}^{3}/8Gt_{\text{ta}}^{2}$. Since any test particle has its own $r_{\text{ta}}$ and $t_{\text{ta}}$, we use the common age of the Universe $t_{U}$ and the common central mass of the object $M$ to express the distance and velocities with respect to $e$ and $\eta$:
$r=r_{0}\frac{1-e\cos\eta}{(\eta-e\sin\eta)^{2/3}}\;,\quad v_{r}=\frac{r_{0}}{t_{U}}\,e\sin\eta\frac{\left(\eta-e\sin\eta\right)^{1/3}}{1-e\cos\eta},$
@@
where $r_{0}$ is related to the mass via: $r_{0}^{3}=GMt_{U}^{2}$.
Figure 1: The normalized velocity-distance relation from the analytical solution (III). The solution is scaled with $r_{0}$ and $r_{0}/t_{U}$, where $t_{U}$ is the age of the Universe. The turnaround $r_{0}$ is related to the enclosed mass. The colored lines corresponds to $e=1$. For different $e$ the area between the colored lines is also covered. The dashed line marks the asymptotic behavior for the unbound solution (which in the case of $H\neq 0$ corresponds to the Hubble flow).
The solutions for different values of $e$ and $\eta$ are illustrated in Fig. 1, which shows the velocity-distance relation for both bounded and unbounded systems and the interplay between the gravity and the local expansion of groups.
@@ IV Expansion Domination
To evaluate the DE’s dominance, specific criteria are utilized. These parameters identify the regimes where dark energy’s repulsive influence overcomes gravitational attraction, governing the evolution of cosmic systems.
@@
IV.1 Hubble Flow around Galaxy Groups and Clusters
Fig. 2 shows an isolated halo from the IllustrisTNG simulations Nelson et al. (2019) with a mass of $\sim 10^{12}\,M_{\odot}$. The upper panel illustrates the bounded and unbounded subhalos around the turnaround. Refs. Peirani and de Freitas Pacheco (2006); Peñarrubia et al. (2014) use different semi-analytical solutions for the local interplay with the Hubble flow, which take the form:
$v(r)=\alpha\,H_{0}\,r-\beta\,\sqrt{\frac{GM}{r}},$ (10)
with $\alpha=1.3$ and $\beta=1.1$. The red line in the figure shows the corresponding fit for the interplay branch. For comparison, we consider observational data on dwarf galaxies in the local environment, taken from McConnachie (2012); McConnachie and Venn (2020a, b); Tully et al. (2023) [^1] LG and Nearby Dwarf Galaxies Link. The lower panel of Fig. 2 shows the radial velocity versus distance from the center of mass of the Local Group (LG). The bounded area, which includes the Milky Way, M31, and their dwarf galaxies, shows a spread up to the turnaround radius. Beyond this, we observe the local interplay to the Hubble expansion.
Asymptotically, the $\ddot{a}/a$ term in Eq. (6) approaches $-\frac{(1+3w)}{2}H^{2}=\Omega_{\Lambda}H_{0}^{2}$, which is the value for a de-Sitter universe. The maximal turnaround, with the condition $\ddot{r}=0$, leads to:
@@ IV.1 Hubble Flow around Galaxy Groups and Clusters
[…]
Figure 3: The curvature phase space for the potential $\Phi$ versus the scalar $\sqrt{\mathcal{K}}$ in a de-Sitter $w=-1$ universe. The Solar System and S-stars around the galactic center are positioned much higher than galaxy groups and clusters, which lie at the transition to local expansion, indicating the dominance of $\Lambda$ in larger structures.
@@
In the context of curvature scalars, the maximal turnaround modifies the Kretschmann scalar to: $\left[\mathcal{K}\right]_{\text{ta}}=3\left[\mathcal{K}\right]_{\text{Cos}}$, where the cosmological Kretschmann scalar is $\left[\mathcal{K}\right]_{\text{Cos}}\equiv 12/D_{H}^{2}$. Here, spacetime curvature is dominated by the background expansion, as the contribution from concentrated mass becomes negligible. If the curvature nears the cosmological value, the expansion dominates motion, as seen for dwarf galaxies outside the turnaround. For the bounded region, the Kretschmann scalar is much higher and Newtonian-dominated, indicating negligible $\Lambda$ effects.
To study the dynamics of galaxy groups and clusters, we analyze systems such as the CenA/M83, LG, M81 groups, and the Virgo, Fornax, and Coma clusters Karachentsev et al. (2013); Tully et al. (2023). We employ their turnaround radii and total enclosed masses to compute curvature parameters. Galaxy groups typically have masses of approximately $\sim 10^{12}\,M_{\odot}$ or greater, while clusters reside in the $\sim 10^{14}\,M_{\odot}$ or higher mass range. As illustrated in Fig. 3, all analyzed systems align with the $\left[\mathcal{K}\right]_{\text{ta}}$ line, demonstrating that dwarf galaxies near the turnaround radius exist at a transition zone between large-scale cosmic expansion and localized gravitational collapse.
@@ IV.2 Spherical Density
The curvature scalar for spherical density $\rho_{m}$ in an expanding $\rho_{\Lambda}$ reads:
@@
where parameter $\kappa$ (from Eq. 4) changes into $\Delta_{c}$ is the ratio between the densities: $\Delta_{c}\equiv\rho_{m}/\rho_{\Lambda}$, with $\rho_{\Lambda}=8\pi G/\Lambda c^{2}$. The parameter emerges in the Kretschmann scalar and quantifies the interplay between the local density of the object and the global cosmological density, that highlights the connection between local gravitational effects and the large-scale influence of dark energy.
This transition is formalized in the Spherical Collapse Model Gunn and Gott (1972); Peebles (1980), which outlines the growth of a small initial density perturbation. The overdensity gradually detaches from the Hubble flow, expands to a maximum turnaround radius and begins to collapse. Collisionless dynamics arrest the collapse, leading to virialization at a final radius which is half of the turnaround radius. The critical overdensity $\Delta_{c}$ quantifies the contrast between the virialized density and the background density at collapse time in a matter-dominated universe. This yields:
$\Delta_{c}\approx 18\pi^{2}\approx 178,$ (14)
implying virialized halos are approximately 178 times denser than their cosmological surroundings or $\Delta_{c}\approx 200$. $\Lambda$CDM cosmology predicts marginally smaller values owing to dark energy’s influence Pace et al. (2017).
@@ IV.3 Binary Motion
In the context of binary motion Benisty et al. (2023, 2024), the curvature scalars for binary systems can be expressed as:
@@
where $\omega_{\text{kep}}$ is the Keplerian orbital frequency. The ratio $\kappa$ naturally emerges from $\mathcal{K}$, highlighting the interplay between orbital dynamics and dark energy:
$\kappa=\left(\frac{T_{\text{Kep}}}{T_{H}}\right)^{2},$ (16)
where $T_{\text{Kep}}=2\pi\sqrt{r^{3}/GM}$ is the Keplerian orbital period, and $T_{H}\equiv 2\pi/H_{0}\approx 27.6\,\text{Gyr}$ is the Hubble time, calculated from the measured Planck values (Aghanim et al., 2020a). To probe the effect of $\Lambda$, one must identify systems with orbital periods close to $T_{H}$. Notably, $T_{H}$ is much larger than the age of the Universe ($\sim 13.7\,\text{Gyr}$), indicating that dark energy’s influence on binary systems is subtle and becomes significant only over cosmological timescales.
The LG has the longest orbital period among bounded binary systems discussed in this paper. Since the orbital period of the LG is about $\approx 0.62T_{H}$, the upper bound on $\Lambda$ is expected to be tighter than those derived from the solar system or the S2 star. Ref. Benisty et al. (2024) finds the upper bound on $\Lambda$ to be 5 times larger than the Planck value Aghanim et al. (2020b). Ref. Benisty et al. (2024) shows that two main characteristics determine the effect of $\Lambda$: the accuracy of the measurements and the period of the system. Although the solar system provides more accurate measurements than the LG, the upper limit on $\Lambda$ derived from the solar system is about $10^{-46}\,\text{m}^{-2}$, which is approximately 10 orders of magnitude larger than the actual value. In contrast, the LG’s longer orbital period results in a tighter upper limit, demonstrating the importance of system timescales in constraining dark energy.
V Curvature Phase Space and Discussion

### v2

@@ III.2 Hubble Flow Around Virilized Structures
We can solve the motion analytically only for the Newtonian case, The canonical Lemaitre–Tolman model describes the dynamics of a spherically symmetric dust cloud under the influence of gravity, without the presence of a cosmological constant ($\Omega_{\Lambda}=0$). The model provides analytical solutions for both bounded and unbounded systems, characterized by the eccentricity parameter $e$ and the phase parameter $\eta$. These solutions describe the evolution of the radial coordinate $r$, the radial velocity $v_{r}$, and the tangential velocity $v_{t}$ as functions of time $t$. For bounded systems, where $e\in[0,1]$ the radial coordinate $r$ and time $t$ are given by Karachentsev et al. (2009):
$r=\frac{r_{\text{ta}}}{2}\left(1-e\cos\eta\right)\;,\quad t=\frac{t_{\text{ta}}}{\pi}\left(\eta-e\sin\eta\right)\;,$ (12)
@@
where $r_{\text{ta}}$ and $t_{\text{ta}}$ are the turnaround radius and turnaround time, respectively. The radial and tangential velocities are:
$v_{r}=\frac{\pi}{2}\frac{r_{\text{ta}}}{t_{\text{ta}}}\frac{e\sin\eta}{1-e\cos\eta}\;,\quad v_{t}=\frac{\pi}{2}\frac{r_{\text{ta}}}{t_{\text{ta}}}\frac{\sqrt{1-e^{2}}}{1-e\cos\eta}\;.$ (13)
These equations describe the oscillatory motion of the system, where the radial coordinate reaches its maximum value at $\eta=\pi$. For the unbounded dwarf galaxies, the radial coordinate $r$ and time $t$ are expressed as:
$r=\frac{r_{\text{ta}}}{2}\left(e\cosh\eta-1\right)\;,\quad t=\frac{t_{\text{ta}}}{\pi}\left(e\sinh\eta-\eta\right)\;,$ (14)
with the corresponding velocities:
@@
This solution describes a system that expands indefinitely, with the radial coordinate growing without bound as $\eta$ increases. The enclosed mass $M_{\text{TA}}$ is given by:
$M_{\text{TA}}=\frac{\pi^{2}r_{\text{ta}}^{3}}{8Gt_{\text{ta}}^{2}},$ (16)
where $t_{U}$ is the age of the Universe. This expression is essentially a manifestation of Kepler’s third law, relating the mass of the system to its dynamical timescale. For bounded systems, $\epsilon$ is negative, indicating a gravitationally bound state:
$r=r_{0}\frac{1-e\cos\eta}{(\eta-e\sin\eta)^{2/3}}\;,\quad v_{r}=\frac{r_{0}}{t_{U}}\frac{e\sin\eta\left(\eta-e\sin\eta\right)^{1/3}}{1-e\cos\eta},$ (17)
and the unbound solution:
@@
Continuing from Fig. (1), the dashed and dot-dashed lines within the bounded region $e\in[0,1]$ illustrate that dwarf galaxies experiencing stronger gravitational encounters (i.e., higher $\eta$ values) exhibit velocities corresponding to smaller radii. This trend indicates that such galaxies occupy dynamically tighter orbits due to energy dissipation from close interactions, confining them to the bound phase of the solution. In contrast, unbound trajectories ($e>1$) at larger $r$ (shaded regions) reflect systems with insufficient binding energy for stable orbits. The concentration of dwarf galaxies at low $r$ underscores the role of environmental effects in shaping their kinematic profiles, aligning with expectations of Newtonian dynamics in high-density regimes.
@@ IV Expansion Domination
To evaluate the DE’s dominance, specific criteria are utilized. These parameters identify the regimes where DE’s repulsive influence overcomes gravitational attraction, governing the evolution of cosmic systems.
IV.1 Hubble Flow around Galaxy Groups and Clusters
System $r_{ta}$ [Mpc] $M_{vir}[M_{\odot}]$ Ref taken Others
Virgo $6.76\pm 0.35$ $(5.7\pm 1.5)\times 10^{14}$ Kim et al. (2020) Sorce et al. (2016); Karachentsev and Nasonova (2010)
Coma $6.67\pm 0.08$ $1.04,2.77]\times 10^{15}$ Benisty et al. (2025)
LG $0.96\pm 0.03$ $\left(2.47\pm 0.15\right)\times 10^{12}$ Makarov et al. (2025) Karachentsev et al. (2009); Peñarrubia et al. (2014)
Fornax [3.88, 5.60][3.88,\,5.60] $1.30,\,3.93]\times 10^{14}$ Nasonova et al. (2011)
M81/M82 $0.89\pm 0.05$ $(2.32\pm 0.31)\times 10^{12}$ Karachentsev et al. (2002)
CenA/M83 $1.55\pm 0.13$ $(6.4\pm 1.8)\times 10^{12}$ Karachentsev et al. (2007)
Table 1: The selected values for the fitted turnaround and enclosed mass from different references.
The velocity distance relation for the Hubble flow line fit assumes the zero peculiar velocity at the big bang $v_{\text{pec}}=0$ and initiate different test particles with initial radial velocities for today under the influence of effectively one body Peirani and de Freitas Pacheco (2006); Peñarrubia et al. (2014) and extended in Del Popolo and Chan (2022). Ref. Peirani and de Freitas Pacheco (2006); Peñarrubia et al. (2014) use different semi-analytical solutions for the local interplay with the Hubble flow. Ref.  Peirani and de Freitas Pacheco (2006) suggest the relation:
$v(r)=1.377H_{0}r-0.976\frac{H_{0}}{r^{n}}\left(\frac{GM}{H_{0}^{2}}\right)^{(n+1)/3},$ (19)
with $n=0.96$. Ref. Peñarrubia et al. (2014) suggest a similar relation for a binary motion:
$v(r)=\left(1.2+0.31\Omega_{\Lambda}\right)\,H_{0}\,r-1.1\,\sqrt{\frac{GM}{r}},$ (20)
where $\Omega_{\Lambda}$ is the DE rate. The red line in the figure shows the corresponding fit for the interplay branch.
@@ IV.1 Hubble Flow around Galaxy Groups and Clusters
Figure 3: The enclosed mass versus turnaround radius (following Pavlidou and Tomaras (2014)) for galaxy groups and clusters. The data points are taken from the references listed in Table under Fig. 4, including Kim et al. (2020); Sorce et al. (2016); Benisty et al. (2025); Makarov et al. (2025); Karachentsev et al. (2009); Peñarrubia et al. (2014); Nasonova et al. (2011); Karachentsev et al. (2002, 2007); Del Popolo et al. (2021). The red curve indicates the Zero Gravity Surface, defined by the balance condition $r_{\Lambda}$ which separates gravitationally bound systems from those dominated by cosmic acceleration.
For comparison, we consider observational data on dwarf galaxies in the LG of galaxies taken from McConnachie (2012); McConnachie and Venn (2020a, b); Tully et al. (2023) [^1] LG and Nearby Dwarf Galaxies Link. We transform the velocities and the distances into the LG center of mass frame. Using the angular separation on the sky $\theta$ the physical distance $r_{\rm{gc}}$ between the galaxy pair is determined by:
$r_{\rm{gc}}^{2}=r_{\rm{g}}^{2}+r_{\rm{c}}^{2}-2r_{\rm{c}}r_{\rm{g}}\cos\theta,$ (21)
where $r_{\rm{g}}$ and $r_{\rm{c}}$ are the distances of the galaxies from the observer. We transform the observed velocities and positions to the center-of-mass frame of the LG. Since the velocity components perpendicular to the line of sight are challenging to observe, the analysis is based only on the observed line-of-sight components. To calculate the radial infall velocities of the galaxies onto the LG center, assumptions about the system dynamics need to be added to estimate the impact of the unobserved velocity components Karachentsev and Kashibadze (2006); Wagner and Benisty (2025). We consider the major infall model that fits for a gravitationally bound systems. The model approximate the radial infall velocity by projecting the velocity difference between a galaxy and the centre of mass onto the line of sight of the galaxy to obtain:
$v_{\mathrm{r,maj}}=\frac{v_{j}-v_{\mathrm{c}}\cos\theta_{\mathrm{c},j}}{{r}_{j}-{r}_{\mathrm{c}}\cos\theta_{\mathrm{c},j}}\;r_{gc}.$ (22)
It thereby assumes that that $v_{\perp,\mathrm{c}}=v_{\mathrm{t}}=0$, in which $v_{\mathrm{t}}$ denotes the tangential velocity. The motion of galaxies within the LG can be described by modeling their velocity field, which provides insights into the system’s dynamics. The analysis begins by correcting observed heliocentric velocities to a Galactocentric frame Karachentsev and Kashibadze (2006); Karachentsev et al. (2009). The LG barycenter is assumed to lie along the line connecting the MW and M31. Its position is determined by their mass ratio $\gamma=M_{MW}/M_{Tot}$. Assuming minimal spread for the velocity dispersion around the center of mass, we get the value $\gamma=0.55$ which is consistent with earlier estimations Karachentsev et al. (2009); Makarov et al. (2025). The right panel of Fig. (2) shows the radial velocity vs. distance from the center of mass of the LG. The bounded area, which includes the MW, M31, and their dwarf galaxies, shows a spread up to the turnaround radius. Beyond this, we observe the local interaction with the local Hubble expansion.
Figure 4: The curvature phase space for the potential $\Phi$ versus the scalar $\sqrt{\mathcal{K}}$ in a de-Sitter w=1w=-1 universe. The Solar System and S-stars around the galactic center are positioned much higher than galaxy groups and clusters, which lie at the transition to local expansion, indicating the dominance of $\Lambda$ in larger structures. The table below gives the characteristic turnaround radii $r_{ta}$ and masses ($M$) of galaxy groups and clusters. Another Ref. for the turnaround and enclosed mass is taken from Kim et al. (2020); Sorce et al. (2016); Benisty et al. (2025); Makarov et al. (2025); Karachentsev et al. (2009); Peñarrubia et al. (2014); Nasonova et al. (2011); Karachentsev et al. (2002, 2007); Del Popolo et al. (2021)
The maximal turnaround is also related to expansion rate, as found by Pavlidou et al. (2014). Asymptotically, the term $\ddot{a}/a$ in Eq. (11) approaches
$\lim_{t\rightarrow\infty}\frac{\ddot{a}}{a}=\Omega_{\Lambda}H_{0}^{2},$ (23)
which is the value for a de Sitter universe. This result highlights that in the far future, the dynamics of the expansion is dominated entirely by the cosmological constant, erasing the influence of matter and radiation. Consequently, any bound system is eventually subject to the repulsive effect of DE, setting a natural limit to the largest scale on which gravity can counterbalance the accelerated expansion. Put differently, there exists a maximum radius beyond which the gravitational attraction of a mass $M$ can no longer hold a test particle against the outward push of DE. At this critical radius, the two effects are exactly in balance. The maximal turnaround obeys the condition $\ddot{r}=0$ leading to:
@@
This distance represents the maximal limit for a test particle within this potential. Ref. Pavlidou et al. (2014) discusses the systems and their turnaround radii, showing that, for bounded galaxy groups and clusters, this is the maximal turnaround. For example, the LG has a maximum turnaround of $1.5\,\text{Mpc}$ (from Eq. (24) with $M=3\cdot 10^{12}\,M_{\odot}$), while Peirani and de Freitas Pacheco (2006); Peñarrubia et al. (2014) reports $r_{\text{ta}}\approx 1\,\text{Mpc}$.
@@
In the context of curvature scalars, the maximal turnaround modifies the Kretschmann scalar to:
$\left[\mathcal{K}\right]_{\text{ta}}=3\left[\mathcal{K}\right]_{\text{Cos}},$ (26)
where the cosmological Kretschmann scalar is:
$\left[\mathcal{K}\right]_{\text{Cos}}\equiv 12/D_{H}^{2}.$ (27)
Here, spacetime curvature is dominated by the background expansion, as the contribution from concentrated mass becomes negligible. If the curvature nears the cosmological value, the expansion dominates motion, as seen for dwarf galaxies outside the turnaround. For the bounded region, the Kretschmann scalar is much higher and Newtonian dominated, indicating negligible $\Lambda$ effects.
To study the dynamics of galaxy groups and clusters, we analyze systems such as the CenA/M83, LG, M81/M82 groups, and the Virgo, Fornax, and Coma clusters Karachentsev et al. (2013); Tully et al. (2023); Kim et al. (2020); Nasonova et al. (2011); Makarov et al. (2025); Karachentsev et al. (2007); Peirani and de Freitas Pacheco (2006); Peirani and Pacheco (2008); Del Popolo et al. (2021); Del Popolo and Chan (2022). For each system, we adopt the characteristic turnaround radii $r_{ta}$ and total enclosed masses $M_{\rm vir}$ summarized in table  (1) and in Fig. (3). Galaxy groups typically have masses of $\sim 10^{12}\,M_{\odot}$, while clusters occupy the $\sim 10^{14}\,M_{\odot}$ or higher mass range. Using these parameters, we compute the corresponding curvature scalars, allowing a direct comparison between local gravitational binding and the influence of cosmic expansion. All of the bounded structures are under the zero-gravity line.
The curvature parameters gives another phase space to compare. Since some curvature scalar vanishes at different cases, we compare $\Phi$ and the $\mathcal{K}$, as Baker et al. (2015). As illustrated in Fig. (4), all analyzed systems align closely with the $\left[\mathcal{K}\right]_{ta}$ line. This indicates that galaxies near the turnaround radius occupy a transitional regime between the dominance of the Hubble flow and the onset of local gravitational collapse, highlighting how DE begins to influence the dynamics of larger structures while smaller, tightly bound systems remain largely decoupled from cosmic expansion.
@@ IV.2 Spherical Density
Another case that our formalism can be formulated with is the spherical density case. The curvature scalar for spherical density $\rho_{m}$ in an expanding $\rho_{\Lambda}$ reads:
@@
where parameter $\kappa$ (from Eq. 7) changes is the known $\Delta_{c}$, which is the ratio between the densities:
$\kappa=\Delta_{c}\equiv\frac{\rho_{m}}{\rho_{\Lambda}},$ (29)
with $\rho_{\Lambda}\equiv 8\pi G/\Lambda c^{2}$. The parameter emerges in the Kretschmann scalar and quantifies the interplay between the local density of the object and the global cosmological density. This highlights the connection between local gravitational effects and the large-scale influence of DE.
[…]
To determine the impact of the local expansion for galaxy groups and clusters we have to take into account objects nearly the turnaround. However, from the spherical collapse picture it gives characteristics for the extension of the bound region of a density. The spherical density framework demonstrates how curvature scalars incorporate the competing influences of local matter density and the uniform background set by DE. This same reasoning can be extended from single, spherically symmetric overdensities to dynamical multi-body systems. In particular, binary motion provides another context in which the interplay between local gravitational binding and cosmic expansion manifests itself in the curvature invariants.
@@ IV.3 Binary Motion
The last case we investigate under the curvature perspective is in the context of binary motion. In the case of binary objects, the curvature scalars for binary systems can be expressed as Benisty et al. (2023, 2024); Strigari (2025),:
@@
where $\omega_{\text{kep}}$ is the Keplerian orbital frequency. The ratio $\kappa$ naturally emerges from $\mathcal{K}$, highlighting the interplay between orbital dynamics and DE:
$\kappa\equiv\left(\frac{T_{\text{Kep}}}{T_{H}}\right)^{2},$ (31)
where $T_{\text{Kep}}=2\pi\sqrt{r^{3}/GM}$ is the Keplerian orbital period, and
$T_{H}\equiv\frac{2\pi}{H_{0}}\approx 27.6\,\text{Gyr}$ (32)
is the Hubble time, calculated from the measured Planck values (Aghanim et al., 2020b). Whereas the spherical density analysis emphasizes virialization and collapse in overdense regions, the binary framework emphasizes the long-term stability of bound orbits against the slow influence of cosmic acceleration. To probe the effect of $\Lambda$, one must identify systems with orbital periods close to $T_{H}$. Notably, $T_{H}$ is much larger than the age of the Universe ($\sim 13.7\,\text{Gyr}$), indicating that DE’s influence on binary systems is subtle and becomes significant only over cosmological timescales.
The LG provides the longest orbital period among bounded binary systems discussed in Benisty et al. (2023). Since the orbital period of the LG is about $\approx 0.62T_{H}$, the upper bound on $\Lambda$ is expected to be tighter than those derived from the solar system or the S2 star. Ref. Benisty et al. (2024) finds the upper bound on $\Lambda$ to be 5 times larger than the Planck value Aghanim et al. (2020a). Ref. Benisty et al. (2024) further shows that two main characteristics determine the effect of $\Lambda$: the accuracy of the measurements and the period of the system. Although the solar system provides more accurate measurements than the LG, the upper limit on $\Lambda$ derived from the solar system is about $10^{-46}\,\text{m}^{-2}$, which is approximately 10 orders of magnitude larger than the actual value. In contrast, the LG’s longer orbital period results in a tighter upper limit, demonstrating the importance of system timescales in constraining DE.
V Discussion